# EconoJax: A Fast & Scalable Economic Simulation in Jax

### v1

@@ Results
Figure 2(b) visualizes the productivity and equality during training of 10 million steps (100.000 tax periods). Introducing taxes increases equality among the population, but this comes at the cost of productivity. This is the result of the productivity-equality trade-off – taxes cause some agents to perform less total labor as they no longer experience an increase in utility. Nevertheless, the government has managed to find tax schedules that overall cause minimal decrease in productivity (Figure 2(b)a). As a result, the government manages to improve on its goal compared to a free market, as shown in Figure 3. Note that, because the population and government agents train simultaneously, an increase in equality does not produce a clear decrease in productivity. This is because the population agents are becoming more efficient during training, causing a natural increase in productivity regardless of taxation.
On average, at the end of training, population agents achieve similar levels of utility in both economic systems, as depicted in Figure 4. However, the median utility, which could be considered a more accurate indicator of overall economic happiness, improves notably with the implementation of taxes, reflecting the increase in equality shown in Figure 2(b)b.
@@ 5. Discussion & Future Work
The experiments conducted in this paper highlight emergent real-world economic behavior. However, this behavior is significantly affected by manually set parameters. In particular, the utility function of the agents has a tunable parameter, and the the coins and labor obtained by each action are set manually as well. All these options heavily influence the simulation. As such, in addition to more functionality added to EconoJax, a necessary improvement would be to align these parameters with the real world, for example by learning them from real data. Nevertheless, due to the manually chosen parameters, care must be taken when building and using simulations such as EconoJax. There is a risk that individuals might unintentionally or deliberately present certain outcomes as genuine economic effects, even though the simulation was designed to produce those effects. However, we believe that the current level of realism in EconoJax is not advanced enough to pose a significant ethical threat to society.
As a multi-agent environment, EconoJax may also serve as an environment for research in multi-agent reinforcement learning. In particular, EconoJax can scale to large action spaces and, by increasing the number of resources in the environment, a large amount of diverse behavior may occur. This makes EconoJax a suitable environment for developing new methods aimed at improving varied behavior in multi-agent systems.
Figure 6. At the top, three examples of action distributions of different agents in the same simulation. It shows agent specialize in different tasks, that are beneficial due to their skill level. These three agents gather and sell, buy and craft, or only trade / perform no labor. The bottom figure shows average prices during the simulation, averaged over 15 environments with the shaded area indicating standard deviation. These environments were initialized with more agents (roughly 2x) proficient in gathering resource 1. As we would expect, the average trading price for this resource is significantly lower.
@@ 6. Conclusion
This paper presented EconoJax, a multi-agent economic simulation written in JAX. EconoJax is a simulation environment in which real-world economic behavior naturally occurs without manually programming specific policies and is largely based on the economy built in the AI economist (Zheng et al., 2022). In contrast to the AI economist, EconoJax allows for fast training of large populations. Compared to their RLlib CPU-based implementation, we saw training time improvements of over 750 times. We conducted experiments with population sizes of 100 and found real-world economic behavior occuring such as specialization, the productivity-equality tradeoff, and realistic tax schedules and market prices. EconoJax can provide a stepping stone for the development of more realistic economic simulations and as an interesting testbed for multi-agent reinforcement learning. In addition, we hope that our results inspire researchers to build upon EconoJax for advanced economic modeling.
Acknowledgements.
This work was financially supported by Shell Information Technology International Limited and the Netherlands Enterprise Agency under the grant PPS23-3-03529461.

### v2

@@ Results
Figure 2(c) visualizes the productivity and equality during training of 10 million steps (100.000 tax periods). Introducing taxes increases equality among the population, but this comes at the cost of productivity. This is the result of the productivity-equality trade-off – taxes cause some agents to perform less total labor as they no longer experience an increase in utility. Nevertheless, the government has managed to find tax schedules that overall cause minimal decrease in productivity (Figure 2(c)a). As a result, the government manages to improve on its goal compared to a free market, as shown in Figure 2(c)c. Note that, because the population and government agents train simultaneously, an increase in equality does not produce a clear decrease in productivity. This is because the population agents are becoming more efficient during training, causing a natural increase in productivity regardless of taxation.
On average, at the end of training, population agents achieve similar levels of utility in both economic systems, as depicted in Figure 3. However, the median utility, which could be considered a more accurate indicator of overall economic happiness, improves notably with the implementation of taxes, reflecting the increase in equality shown in Figure 2(c)b.
@@
5. Multi-Agent Methods
(a) Agent Returns (r=4)
(b) Agent Returns (r=12)
Figure 5. Agent returns with various standard multi-agent practices for different amount of resources in EconoJax. Higher resources create a larger action space and more possibilities for agents to produce diverse behavior. Our results indicate that centralized training produces roughly the same behavior compared to training individual networks.
[…]
In this section, we perform additional experiments in EconoJax with an increasing number of resources. By increasing the number of resources, the possible variety in agent behavior grows, which allows us to test the effect of different multi-agent methods in terms of their diversity.
For these experiments, we inspect three common methods for multi-agent RL: 1) centralized training, sharing all network parameters among all agents. Agents will behave differently due to the different private observations given to the network for each agent; 2) Independent training, each agent training their own set of network parameters; 3) In actor-critic methods, where RL agents train a value- and policy network separately, we can mix both prior methods by training individual policy networks and a central value network. Methods utilizing this technique are often referred to as centralized training, decentralized execution (CTDE) methods. For this last method, we only consider the naive version in which the value network observes the same type of observation as the policy network (with larger batches due to the network observing the observation for each agent). Although there are more sophisticated CTDE methods Kuba et al. (2022) that may obtain higher performance, EconoJax agents are tasked with mimicking the real world and enhancing their value network observation arguably degrades their realism.
We used PPO Schulman et al. (2017) as the learning algorithm for all three of our methods. The independent varient is then sometimes referred to as iPPO, and the sharing of the value network as MAPPO Yu et al. (2022). In these experiments, we only consider population agents and disabled taxes. The parameters of the previous experiments shown in Table 2 are reused with a few exceptions: First, we vary the number of resources and set $R_{d}$ to $\text{min}(1,log_{2}(\text{num resources}))$. Next, we initialize all skills with a normal distribution around 1.0 with a standard deviation of 1.0 and allow agents to increase their skill. Whenever an agent gathers or crafts, the respective skill is increased by 0.5%. This 0.5% bonus shrinks as the agent approaches the maximum skill level of 5. Additionally, we slightly simplified the action space by allowing the agents to only place trade orders for three prices: 3, 6 and 9. Lastly, we increase the network size to 256 nodes per layer for all shared networks.
Figure 5(b) showcases the effect of the different methods in EconoJax with 4, and 12 different resources – in which the agents have a total of 30 and 86 discrete actions, respectively. We can see that even in larger action spaces, different learning methods achieve similar returns. Intuitively, however, we expect individually trained policies to be more diverse Johanson et al. (2022), as they only learn from their own observations and rival other agents. In turn, we would assume that this highly influences the outcome of the simulation. Instead, our experiments indicate that centralized learning may well be capable of diverse learning diverse behavior.
Naturally, returns alone could arrive at similar levels via different policies. As such, we additionally measured the spread of the action distributions of each agent, where we binnend action types belonging to specific resource types (e.g. binning all sell resource 1 actions). When averaging these action distributions over multiple evaluations, we again find no noticeable difference between the different multi-agent learning methods.
We note that, in order to produce more heterogeneous policies in centralized trained agents, some works enhance the observations with an agent id Rashid et al. (2020). We included this method in our experiments and again found no noticeable differences in action distributions or final returns as shown in Figure 5(b). In fact, the learning pattern almost exactly mirrors that of centralized training without agent ids.
We believe that, given that the state contains sufficient information for specialized behavior (in the case of EconoJax likely by observing the skill levels), centralized training should be effective, even in large action spaces. Centralized training offers the advantage of reduced computational demands as well as allowing the population to grow dynamically without the need to heavily retrain agents.
@@ 6. Discussion & Future Work
The experiments conducted in this paper highlight emergent real-world economic behavior. However, this behavior is significantly affected by manually set parameters. In particular, the utility function of the agents has a tunable parameter, and the coins and labor obtained by each action are set manually as well. All these options heavily influence the simulation. As such, in addition to more functionality added to EconoJax, a necessary improvement would be to align these parameters with the real world, for example by learning them from real data. Nevertheless, due to the manually chosen parameters, care must be taken when building and using simulations such as EconoJax. There is a risk that individuals might unintentionally or deliberately present certain outcomes as genuine economic effects, even though the simulation was designed to produce those effects. However, we believe that the current level of realism in EconoJax is not advanced enough to pose a significant ethical threat to society.
We experimented with different multi-agent methods and action space sizes and observed that centralized training may provide adequate diverse behavior even in action spaces of size 86. Future work may experiment with even larger experiments and provide a better metric to measure diversity.
Figure 6. Top: three action distributions of different agents in the same simulation, showcasing specialization in different tasks, that are beneficial due to their skill level. These three agents gather and sell, buy and craft, or only trade / perform no labor. The bottom figure shows prices in an environment episode, averaged over 15 environments. These environments were initialized with more agents (roughly 2x) proficient in gathering resource 1. As we would expect, the average trading price for this resource is noticeably lower.
@@ 7. Conclusion
This paper presented EconoJax, a multi-agent economic simulation written in JAX. EconoJax is a simulation environment in which real-world economic behavior naturally occurs without manually programming specific policies and is largely based on the economy built in the AI economist Zheng et al. (2022). In contrast to the AI economist, EconoJax allows for fast training of large populations. Compared to their RLlib CPU-based implementation, we saw training time improvements of over 2.400 times. We conducted experiments with population sizes of 100 and found real-world economic behavior occuring such as specialization, the productivity-equality tradeoff, and realistic tax schedules and market prices. EconoJax can provide a stepping stone for the development of more realistic economic simulations. We further showcased how different multi-agent learning methods did not have a significant effect on the policies produced after training on varying action space sizes.
{acks}
This work was in part performed using the compute resources from the Academic Leiden Interdisciplinary Cluster Environment (ALICE) provided by Leiden University. This work was financially supported by Shell Information Technology International Limited and the Netherlands Enterprise Agency under the grant PPS23-3-03529461. We would lastly like to thank Felix Kleuker for being a mathematician.
Appendix A Environment Differences Compared to the AI Economist
This section aims to summarize implementation differences between the EconoJax environment and the original AI economist environment.
EconoJax agents do not move in a 2D world in which they have to collect resources and build houses. The ‘move‘ action has therefor been replaced by a per-resource ‘gather‘ action which allows an agent to instantly receive a resource. This additionally means that, in the base version of EconoJax, resources are not scarce. However, time is still scarce, forcing agents to specialize. Furthermore, building a house no longer requires a free spot in the world, and agents can not block each other via the building of houses.
Due to the removal of the 2D spatial world, EconoJax agents observe a flat input space, rather than a combined 2D and 1D space. Additionally, the observations of the market have been simplified to relevant statistics (highests/lowests/number of orders) instead of the full market.
The government agent’s observations have also been altered. In the AI Economist, individual states about each agent were observed. In EconoJax, the government observes mean, median, and std statistics for various population metrics. This is largely because EconoJax can comfortably feature a much larger population count, and feeding observations for each would be infeasible and unrealistic from a policy-making standpoint.
EconoJax is less explicit about the type of resources available in the world. Rather than ”Wood” and ”Stone”, there is simply $n$ amount of resources available, and crafting these into coins requires a configurable amount of resources.
Optionally, in EconoJax, agents can improve their skill in crafting (building houses) or gathering resources by performing these actions. The rate at which their skill improves is configurable.
The EconoJax resource market differs slightly in the way the price is determined for a matched trade. In the AI Economist, when two orders are matched, the price is set to whichever order was placed first. In EconoJax, we use the price of the order that was placed last. This is altered because (in both environments) the lowest/highest asks/bids are evaluated for matching before other orders. If the price were to be set by the first order, agents could exploit the market by placing only very low/high orders whenever a matching order already exists.